# Uncertainty quantification for intervals

### v1

@@ Abstract
Data following an interval structure are increasingly prevalent in many scientific applications. In medicine, clinical events are often monitored between two clinical visits, making the exact time of the event unknown and generating outcomes with a range format. As interest in automating healthcare decisions grows, uncertainty quantification via predictive regions becomes essential for developing reliable and trusworthy predictive algorithms. However, the statistical literature currently lacks a general methodology for interval targets, especially when these outcomes are incomplete due to censoring. We propose a uncertainty quantification algorithm and establish its theoretical properties using empirical process arguments based on a newly developed class of functions specifically designed for interval data structures. Although this paper primarily focuses on deriving predictive regions for interval-censored data, the approach can also be applied to other statistical modeling tasks, such as goodness-of-fit assessments. Finally, the applicability of the methods developed here is illustrated through various biomedical applications, including two clinical examples: i) sleep time and its link with cardiovasculuar diseases ii) survival time and physical activity values.
62F,
62G,
62N,
Interval-censoring, Conformal inference, Empirical processes,
keywords:
[class=MSC]
@@
1 Introduction
Survival analysis is a core statistical methodology in biomedical research and precision medicine [40], providing methods to quantify the lifetime of patients under various conditions and treatments. In these fields, survival models are the primary tools for establishing the effectiveness of drugs and clinical interventions concerning the occurrence of specific events, as is the case in clinical trials [18].
Traditionally, practitioners have focused on right-censored outcomes, which constitute the conventional setting in survival analysis [38]. However, in other clinical applications, such as digital health, outcomes often possess an interval structure. For instance, with electronic records [34], the time to a clinical event may occur between two medical visits, as seen in the case of diabetes mellitus or other chronic diseases. Another example of an interval setup is found in clinical settings related to sleep studies. Here, patients typically record the time they go to bed and the time they wake up; however, the precise moment they actually fall asleep is not documented. This results in an interval of time during which the event (falling asleep) occurs, but the exact time within this interval remains unknown. Similarly, with wearables, devices are not worn continuously, and some defined clinical events may occur between two wear periods. Given that patient monitoring is often quasi-continuous, it is imperative to develop new methodologies for handling these interval-structured outcomes. A comprehensive illustration is depicted in Figure 1.
Figure 1: Illustration of a target variable with an interval structure. Each segment represents a patient. The borders of the black lines indicate medical checkpoints that bound the true, unknown disease onset. The light grey segments represent right-censored patients, while the dark grey segments represent left-censored patients.
Quantifying the uncertainty of predictive outcomes is one of the most crucial modeling tasks in statistical sciences, particularly in clinical applications within personalized medicine [30]. In clinical studies, the response of patients over time exhibits individual and heterogeneous behavior. Therefore, when predicting clinical outcomes, there is considerable variability in patient responses. Reporting only point estimates, such as the conditional mean, provides partial and imperfect knowledge of the phenomenon under study and can lead to misleading clinical conclusions regarding the effects of drugs and interventions [5]. Consequently, measures of uncertainty, such as predictive regions, should be included in analyses. Although the construction of covariate-dependent predictive regions has received significant attention throughout the historical development of statistics [70, 69], the literature on survival analysis and censored outcomes remains more limited. With the proliferation of data-driven systems across various scientific and engineering fields, quantifying uncertainty has become a central research topic within the statistical and machine learning communities.
A particularly prolific framework that has emerged in recent years is conformal inference [69], which encompasses a general family of methods. Given a random sample $\mathcal{D}_{N}=\{(X_{i},T_{i})\in\mathcal{X}\times\mathcal{T}:1\leq i\leq N\}$, assumed to be at least exchangeable, and a new observation $(X_{N+1},T_{N+1})\in\mathcal{X}\times\mathcal{T}$ with the same exchangeability property relative to $\mathcal{D}_{N}$, we can define a predictive region $\widehat{\mathcal{C}}_{1-\alpha}(\cdot)\subset\mathcal{T}$ such that the following non-asymptotic property is satisfied:
$\mathbb{P}(T_{N+1}\in\widehat{\mathcal{C}}_{1-\alpha}(X_{N+1}))\geq 1-\alpha.$
Here, $\mathbb{P}$ represents the joint probability law over the entire dataset $\mathcal{D}_{N}$ and the new data point $(X_{N+1},T_{N+1})$. Other methods in the data analysis literature that address similar scientific problems include Bayesian methods, which require the subjective specification of priors for automatic uncertainty quantification, asymptotic approximations based on the central limit theorem, and, more broadly, bootstrapping techniques [61].
The key characteristic of the various methods mentioned is that they operate on scalar, multivariate, and functional random variables $T\in\mathcal{T}$ (see review [49]). However, in several scientific applications, the only available information is a random interval $(L,U)$, where $0\leq L\leq U\leq\infty$ almost surely, and almost surely $T\in(L,U)\subset\mathcal{T}$. This setup encompasses right-censoring when $U=\infty$ and left-censoring when $L=0$.
@@ 1 Introduction
General methodologies for providing predictive regions of an interval-censored target typically focus on estimating the conditional distribution [32], but they may result in poorly calibrated predictive intervals and a higher dependence on the chosen model. The non-parametric estimation of survival functions with interval-censored data is based on [66]’s algorithms, which can be derived using a self-consistency argument [16]. Efron also proved that this approach yields the nonparametric maximum likelihood estimator (NPMLE). However, significant issues highlighted in the literature include the lack of uniqueness when the log-likelihood is not strictly concave, and the fact that consistency results are typically available only under the assumption of a fixed number of inspection times [26].
In [50] the authors deduce that the NPMLE for the joint distribution function of an interval-censored survival time and a continuous mark variable is generally not consistent. Furthermore, the self-consistent approach may not be consistent even when the inspection times are limited to finitely many values, as demonstrated by a counterexample in [76]. Additionally, traditional methodologies in computational statistics, such as the bootstrap, are known to be inconsistent when using the NPMLE as the reference measure in the presence of interval censoring [57]. Due to these challenges, practitioners often resort to parametric and semi-parametric approaches for modeling interval-censored data, as these methods are numerically more stable and yield consistent solutions from both optimization and statistical perspectives [53].
@@
To address this gap in the literature, the goal of this paper is to introduce a new uncertainty quantification algorithm, uncervals. This algorithm is theoretically grounded in a blend of conformal prediction and bootstrap methods [57], following the resampling ideas proposed by [55]. Our approach achieves significantly better calibration, as confirmed by our simulation studies, which are consistent with previous work (see, for example, [78]). The new method could be applied to develop novel algorithms for interval-censored data, including variable selection methods and goodness-of-fit tests. To derive the theoretical properties of uncervals, we introduce a new class of functions that leverage empirical process theory. This new uncertainty interval quantification framework is highly relevant in clinical applications for precision medicine and digital health, where outcomes often exhibit an interval structure.
@@
    Theoretical support: To derive the formal guarantees of uncervals, we introduce a new class of functions that inherently capture the statistical nature of interval structures. This enables the use of empirical process theory to establish consistency and convergence rates.
@@
    Biomedical applications: We illustrate the potential of uncervals through relevant cases involving interval-censored data. We compare our approach to naive adaptations of right-censored algorithms and the consideration of quantiles for conditional survival functions. For these purposes, we examine the performance of our proposal across different scenarios using parametric, semi-parametric, and non-parametric regression algorithms.
@@ 1.1 Paper Structure
The structure of the paper is defined as follows. Section 2 describes the key literature on uncertainty quantification and conformal prediction for right-censored data, providing a natural reference context. Section 3 introduces the algorithms for uncertainty quantification of interval-censored data and their theoretical analysis (Section 3.1). Section 4 presents a simulation analysis to examine the theoretical properties of uncervals. Section 5 introduces two examples to illustrate the potential of the methodology in biomedical applications. Finally, Section 6 discusses the contributions, limitations and future research directions of our interval uncertainty quantification framework.
1.2 Notation
$e$ Operating mode (0: deterministic, 1: randomized)
$\widehat{\mathcal{C}}^{e}_{1-\alpha}(\cdot)$ Predictive region
$\widehat{L}_{1-\alpha}(\cdot)$ Lower Predictive Bound
$T_{i}$ Unobserved underlying time
$L_{i},U_{i}$ Lower and upper bounds of a time interval, respectively
$X_{i}$ p-dimensional covariates
$F$ True conditional distribution function
$\hat{F}_{1}$ Estimator of the conditional distribution function
$\Phi_{i}$ Conditional probability integral transform of $T_{i}$ through $\hat{F}_{1}$
$\Lambda_{i},\Upsilon_{i}$ Evaluations of $\hat{F}_{1}$ on observable data $(L_{i},X_{i})$ and $(U_{i},X_{i})$ respectively
$\mathbb{P}$ Joint probability law over $\mathcal{D}_{N}$ and $(X_{N+1},Y_{N+1})$
$\mathbf{P}$ Marginal probability law of split 1 $\mathbf{P}(\cdot)=\mathbb{P}(\cdot\cap\mathcal{I}_{1})$
${P}$ Probability law conditional on split 1 $P(\cdot)=\mathbb{P}(\cdot|\mathcal{I}_{1})$
$n$ Last index in split 2
$N$ Last index in split 1 and in $\mathcal{D}_{N}$
$G(t)$ Population cdf of $\Phi_{1},\ldots,\Phi_{n}$ conditional on $\hat{F}_{1}$
$\mathbb{Q}_{n}(t)$ Interval measure, $\mathbb{Q}_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\Upsilon_{i}-\Lambda_{i}}1_{(\Lambda_{i},\Upsilon_{i})}(t)$

### v2

@@ Abstract
Data following an interval structure are increasingly prevalent in many scientific applications. In medicine, clinical events are often monitored between two clinical visits, making the exact time of the event unknown and generating outcomes with a range format. As interest in automating healthcare decisions grows, uncertainty quantification via predictive regions becomes essential for developing reliable and trustworthy predictive algorithms. However, the statistical literature currently lacks a general methodology for interval targets, especially when these outcomes are incomplete due to censoring. We propose an uncertainty quantification algorithm for interval responses and establish its theoretical properties using empirical process arguments based on a newly developed class of functions specifically designed for these interval data structures. Although this paper primarily focuses on deriving predictive regions for interval-censored data, the approach can also be applied to other statistical modeling tasks, such as goodness-of-fit assessments. Finally, the applicability of the method is demonstrated through simulations, showing up to a 60% improvement in conditional coverage. Our new algorithm is also applied to various biomedical contexts, including two clinical examples: i) sleep duration and its association with cardiovascular diseases, and ii) survival time in relation to physical activity levels.
Keywords: Interval-censoring, Conformal inference, Empirical processes.
1 Introduction
Survival analysis constitutes core statistical methodology in biomedical research and precision medicine (Kosorok & Laber, 2019), providing methods to quantify the lifetime of patients under different conditions and treatments. In these fields, survival models are the primary tools for establishing the effectiveness of drugs and clinical interventions concerning the occurrence of specific events, as is the case in clinical trials (Fleming & Lin, 2000).
Traditionally, practitioners have focused on right-censored outcomes, which constitute the conventional setting in survival analysis (Kaplan & Meier, 1958). However, in other clinical applications, such as digital health, outcomes often possess an interval structure. For instance, with electronic records (Hyun et al., 2017), the time to a clinical event may occur between two medical visits, as seen in the case of diabetes mellitus or other chronic diseases. Another example of an interval setup is found in clinical settings related to sleep studies. Here, patients typically record the time they go to bed and the time they wake up; however, the precise moment they actually fall asleep is not documented. This results in an interval of time during which the event (falling asleep) occurs, but the exact time within this interval remains unknown. Similarly, with wearables, devices are not worn continuously, and some defined clinical events may occur between two wear periods. Given that patient monitoring is often quasi-continuous, it is imperative to develop new methodologies for handling these interval-structured outcomes. A comprehensive illustration is depicted in Figure 1.
Figure 1: Illustration of a target variable with an interval structure. Each segment represents a patient. The borders of the black lines indicate medical checkpoints that bound the true, unknown time-to-event. The light grey segments represent right-censored patients, while the dark grey segments represent left-censored patients.
Quantifying the uncertainty of predictive outcomes is one of the most crucial modeling tasks in statistical sciences, particularly in clinical applications within personalized medicine. In clinical studies, the response of patients over time exhibits individual and heterogeneous behavior. Therefore, when predicting clinical outcomes, there is considerable variability in patient responses. Reporting only point estimates, such as the conditional mean, provides partial and imperfect knowledge of the phenomenon under study and can lead to misleading clinical conclusions regarding the effects of drugs and interventions (Candès et al., 2023). Consequently, measures of uncertainty, such as predictive regions, must be included in analyses. Although the construction of covariate-dependent predictive regions has received significant attention throughout the historical development of statistics (Wilks, 1941; Vovk et al., 2005), the literature on survival analysis and censored outcomes remains more limited. With the proliferation of data-driven systems across various scientific and engineering fields, quantifying uncertainty has become a central research topic within the statistical and machine learning communities.
A particularly prolific framework that has emerged in recent years is conformal inference (Vovk et al., 2005), which encompasses a general family of methods. Given a random sample $\mathcal{D}_{N}=\{(X_{i},T_{i})\in\mathcal{X}\times\mathcal{T}:1\leq i\leq N\}$, assumed to be at least exchangeable, and a new observation $(X_{N+1},T_{N+1})\in\mathcal{X}\times\mathcal{T}$ with the same exchangeability property relative to $\mathcal{D}_{N}$, it is possible to consider a predictive region $\widehat{\mathcal{C}}_{1-\alpha}(\cdot)\subset\mathcal{T}$ based on $\mathcal{D}_{N}$ such that the following non-asymptotic property is satisfied: $\mathbb{P}(T_{N+1}\in\widehat{\mathcal{C}}_{1-\alpha}(X_{N+1}))\geq 1-\alpha.$
Here, $\mathbb{P}$ represents the joint probability law over the entire dataset $\mathcal{D}_{N}$ and the new data point $(X_{N+1},T_{N+1})$. Other methods in the data analysis literature that address similar scientific problems include Bayesian methods, which require the subjective specification of priors, asymptotic approximations based on the central limit theorem, and, more broadly, bootstrapping techniques (Stine, 1985).
@@ 1 Introduction
General methodologies for providing predictive regions of an interval-censored target typically focus on estimating the conditional distribution (Huang, 1996), but they may result in poorly calibrated predictive intervals and a higher dependence on the chosen model. The non-parametric estimation of survival functions with interval-censored data is based on (Turnbull, 1976)’s algorithms, which can be derived using a self-consistency argument (Efron, 1967). Efron also proved that this approach yields the nonparametric maximum likelihood estimator (NPMLE). However, significant issues highlighted in the literature include the lack of uniqueness when the log-likelihood is not strictly concave, and the fact that consistency results are typically available only under the assumption of a fixed number of inspection times (Gentleman & Geyer, 1994). (Maathuis & Wellner, 2008) deduce that the NPMLE for the joint distribution function of an interval-censored survival time and a continuous mark variable is generally not consistent. Furthermore, the self-consistent approach may not be consistent even when the inspection times are limited to finitely many values, as demonstrated by a counterexample in (Yu et al., 2000). Additionally, traditional methodologies in computational statistics, such as the bootstrap, are known to be inconsistent when using the NPMLE as the reference measure in the presence of interval censoring (Sen & Xu, 2015). Due to these challenges, practitioners often resort to parametric and semi-parametric approaches for modeling interval-censored data, as these methods are numerically more stable and yield consistent solutions from both optimization and statistical perspectives (Meixide et al., 2024).
To address this gap in the literature, the goal of this paper is to introduce a new uncertainty quantification algorithm, uncervals. Our proposal is inspired by a combination of conformal prediction and bootstrap methods (Sen & Xu, 2015), following the results in Politis (2015), which suggest that bootstrap improves conditional coverage. This serves as the rationale for combining conformal prediction with bootstrap in our approach: to enhance finite-sample performance in terms of marginal and conditional coverage. See the simulations in Section 3.1.
@@
Our approach achieves significantly better calibration, as confirmed by our simulation studies, which are consistent with previous work (see, for example, Zhang (2022)). The new method could be applied to develop novel algorithms for interval-censored data, including variable selection methods and goodness-of-fit tests. To derive the theoretical properties of uncervals, we introduce a new class of functions that leverages empirical process theory and turns out to be highly relevant in clinical applications for precision medicine and digital health, where outcomes often exhibit an interval structure.
@@
    Theoretical support: To derive the formal guarantees of uncervals, we introduce a new class of functions that inherently capture the statistical nature of interval data. This enables the use of empirical process theory to establish consistency and convergence rates.
@@
    Biomedical applications: We illustrate the potential of uncervals through relevant cases involving interval-censored data. We compare our approach to naive interval-censored adaptations of algorithms originally crafted for right-censoring. For these purposes, we examine the performance of our proposal across different scenarios and semi-parametric models.
@@ 1.1 Paper Structure
The structure of the paper is defined as follows. Section 1.2 describes the key literature on uncertainty quantification and conformal prediction for right-censored data, providing a natural reference context to depart to interval censoring. Section 2 introduces our algorithms for uncertainty quantification with interval-censored data and their theoretical analysis (Section 2.2). Section 3 presents a simulation analysis to examine the theoretical properties of uncervals. Section 4 introduces two examples to illustrate the potential of the methodology in biomedical applications. Finally, Section 5 discusses the contributions, limitations and future research directions of our framework.
1.2 Related work
We present a literature review to derive prediction regions in survival analysis based on conformal prediction techniques. Notably, our proposal is designed for interval-censored data, whereas the existing literature focuses on right-censored outcomes. The right-censored case can be considered a special case of interval-censored data; however, despite this, our method transcends being a mere extension of right-censored approaches, which do not have the potential to operate effectively for interval-censored data.
Survival analysis traditionally aims to infer the probability that a patient will survive beyond a specified time. This is often complicated by the presence of censoring, where for some patients, survival times are only known to exceed a certain value due to limitations such as the end of a study period. Machine learning methods are introduced as promising tools for handling such complex data without relying on strong modeling assumptions. However, these methods face challenges in quantifying uncertainty, which is critical for making reliable predictions in high-stakes situations. To overcome this problem, Candès et al. (2023) present a methodology to construct prediction intervals for survival times under the Type-I censoring setting, which assumes that the censoring time is observed for every unit, while the outcome is only observed for uncensored units. This is more restrictive than traditional right-censoring, where the censoring time is not available for censored units. This new lens on the problem is particularly relevant in modern healthcare applications, such as the COVID-19 pandemic, where accurate prediction of survival times is crucial for resource allocation and decision-making in public health crises. Refer to the supplement for a comprehensive review of the statistical literature on conformal prediction.
The type of predictive region Candès et al. (2023) propose is called the lower predictive bound (LPB), which can be seen as a one-sided tolerance region. The LPB serves as a conservative estimate of the survival time, offering a critical tool for high-stakes decision-making by ensuring that the true survival time falls above this bound with a specified probability. We say an LPB $\hat{L}_{1-\alpha}(\cdot)$ is calibrated if $\mathbb{P}(T_{N+1}\geq\widehat{L}_{1-\alpha}(X_{N+1}))\geq 1-\alpha,$ where $\alpha$ is a pre-specified level, and the probability is computed over both $\mathcal{D}_{N}$ and a future unit $(X_{N+1},T_{N+1})$ that is independent of $\mathcal{D}_{N}$. This can be seen as a predictive interval whose upper border is +$\infty$+ ∞, serving as a conservative assessment of survival.
Let $\tilde{T}_{i}=\min(T_{i},C_{i})$. Upon making the following elementary observation:
$\tilde{T}_{i}\leq T_{i}\Longrightarrow\textrm{ any calibrated LPB on }\tilde{T}_{i}\textrm{ is also a calibrated LPB on }T_{i}$ (1)
a naive approach to construct distribution-free LPBs in right-censored scenarios immediately arises: just apply conformal prediction over the censored target $\tilde{T}$ in a one-sided fashion. However, this approach is overconservative in the sense that it will lead to an excessively low $\hat{L}_{1-\alpha}(\cdot)$. Candès et al. (2023) observe that this phenomenon is related to censoring times being overall smaller than survival times. This led them to apply conformal inference on subpopulations with larger censoring times while correcting for the covariate shift this selection produces applying a one-sided version of weighted conformal inference (Tibshirani et al., 2019). It can even be shown that if $(T,X)\perp C$, then the covariate shift vanishes and the technique reduces to applying conformal inference on $\left(X_{i},T_{i}\wedge c_{0}\right)_{C_{i}\geq c_{0}}$. This has two main problems:
  • In general, there is no access to all the censoring times in practical situations. This would involve, for instance, knowing when a person would quit from a clinical trial even though this individual has experienced the event of death.
  • Their approach depends on a hyperparameter $c_{0}$ that selects a subpopulation with censoring times bigger than $c_{0}$. Figure 2(a) serves as an illustration of how sensitive the approach by Candès et al. (2023) is to the hyperparameter $c_{0}$.
(a) Evolution of empirical coverage with the threshold $c_{0}$ used in Candès et al. (2023). The desirable outcome would be a horizontal line at 0.9. Setup details in 3.4.
(b) Red: conditional coverage achieved by naively taking the $\alpha$-th quantile of a distributional regression algorithm. Black: our approach. Details in Section 3.1
Figure 2: Failure of uncertainty quantification approaches for interval data
Generalizing the existing non-asymptotic guarantees for right-censored data to new ones that account for interval-censoring is highly non-trivial. This difficulty arises in part from the fact that, under interval-censoring, survival times are never observed—we only have access to $(L,U)$, which almost surely contains the true survival time $T$. We begin our arguments by building upon an observation analogous to (1):
$L_{i}\leq T_{i}\Longrightarrow\textrm{ any calibrated LPB on }L_{i}\textrm{ isalso a calibrated LPB on }T_{i}$ (2)
The first immediate issue of this approach is that it would lead to even less tight LPBs. However, there is a clear benefit from this, which is the well-known finite sample size coverage guarantee inherent to conformal inference, as stated in Theorem 1.
An interval whose lower and upper borders are respectively the $\frac{\alpha}{2}$ and $1-\frac{\alpha}{2}$ quantiles of $T$ given $X=x$ is conditionally valid. This would be brought into practice by plugging consistent estimators of such quantiles, recoverable from the conditional survival function estimated by, for instance, interval-censored recursive forests (Hunyong Cho & Kosorok, 2022).
Other recent notable work include Sesia & Svetnik (2024), which proposes a conformal inference method for constructing lower prediction bounds for survival times from right-censored data by imputing unobserved censoring times and then accounting for covariate shift (Tibshirani et al., 2019). Farina et al. (2025) introduce two new methods for right-censored outcomes based on inverse-probability-of-censoring weighting and augmented inverse-probability-of-censoring weighting.
These ideas, originally designed for right-censoring, do not easily transfer to the context of interval-censored outcomes. This due to the fact that in right-censored we actually manage to observe times, while in interval-censoring point observations are never sampled. Our strategy differs both theoretically and practically, as we will see next.
2 Predictive regions for interval-censored targets
2.1 Setup and notation
[…]
Let $F(t,x)=P\left(T_{i}\leq t\mid X_{i}=x\right)$ denote the conditional cumulative distribution function (cdf) of $T_{i}$ given $X_{i}=x$. Throughout the paper, we will split the whole data $\{1,\ldots,N\}$ into $\mathcal{I}_{1}:=\{n+1,\ldots,N\}$ and $\mathcal{I}_{2}:=\{1,\ldots,n\}$. Because of being our data i.i.d, random splits are also possible. We assume we have obtained an estimator $\hat{F}_{1}$ of $F$ based on $\mathcal{I}_{1}$ and we denote by $\Lambda_{i}:=\hat{F}_{1}(L_{i},X_{i})$, $\Upsilon_{i}:=\hat{F}_{1}(U_{i},X_{i})$ the evaluations of $\hat{F}_{1}$ on observable data coming from $\mathcal{I}_{2}$ and indexed by $i=1,\ldots,n$ according to how it was chosen. We denote by $\Phi_{i}:=\hat{F}_{1}(T_{i},X_{i})$ the evaluations of $\hat{F}_{1}$ on the underlying survival times $T_{i}$ (we do not have empirical access to $\Phi_{i}$, but we need to postulate them). Upon fixing $\left\{(L_{i},U_{i},X_{i})\right\}_{i\in\mathcal{I}_{1}}$- and therefore also $\hat{F}_{1}(\cdot,\cdot)$- then $\left\{(\Phi_{i},\Lambda_{i},\Upsilon_{i})\right\}_{i\in\mathcal{I}_{2}}$ are i.i.d. because they are evaluations on an i.i.d. sample of a function that became deterministic after conditioning. We denote by ${P}$ the conditional-on-split-1 joint law of $(\Phi,\Lambda,\Upsilon)$ and by $G$ the population cdf of $\Phi_{1},\ldots,\Phi_{n}$ conditional on $\hat{F}_{1}$, respectively. $\mathbb{P}$ denotes the joint probability law over $\mathcal{D}_{N}$ and $(X_{N+1},T_{N+1})$; $\mathbf{P}(\cdot)=\mathbb{P}(\cdot\cap\mathcal{I}_{1})$ the marginal probability law of split 1; and $P(\cdot)=\mathbb{P}(\cdot|\mathcal{I}_{1})$ the probability law conditional on split 1. The sample structure is as follows, where we chose split 2 to be before split 1 just for notational convencience:
$\underbrace{1,\ldots,n-1,n}_{\text{split 2}},\underbrace{n+1,\ldots,N-1,N}_{\text{split 1}},\underbrace{N+1}_{\text{new}}.$
We present uncervals in Algorithm 1, where the user is free to choose between two operating modes indexed by $e=0,*$.
Algorithm 1 uncervals
1:$\left\{(L_{i},U_{i},X_{i})\right\}_{i=1}^{N}$, $\alpha$, $X_{N+1},e$
2:
4:Obtain $\hat{F}_{1}$ ($1-$ conditional survival function) based on $\mathcal{I}_{1}$ using an off-the-shelf semi-parametric model for interval-censored responses.
5:for $i=1,\ldots,n$ do
6:     $j\leftarrow\textrm{Uniform}(\{1,\ldots,n\})$
7:     if $e=0$ then
8:         $\Phi^{*}_{i}\leftarrow$ $\hat{F}_{1}(L_{j},X_{j})$
9:     else
10:         $\Phi^{*}_{i}\leftarrow$ $\hat{F}_{1}(L_{j},X_{j})+(\hat{F}_{1}(U_{j},X_{j})-\hat{F}_{1}(L_{j},X_{j}))\cdot\textrm{Uniform}(0,1)$
11:     end if
12:     $V_{i}^{*}\leftarrow\psi\left({\Phi}_{i}^{*}\right)$
13:end for
14:$\hat{Q}_{\mathcal{I}_{2}}^{*}\leftarrow(1-\alpha)\left(1+1/n\right)$ empirical quantile of $\left\{V_{i}^{*}\right\}$ Final $(1-\alpha)$ prediction set $\widehat{\mathcal{C}}_{1-\alpha}^{e}(X_{N+1})\leftarrow$$\left\{t\geq 0:\psi(\hat{F}_{1}(t,X_{N+1}))\leq\hat{Q}^{*}_{\mathcal{I}_{2}}\right\}$
In Algorithm 1, $\psi(\cdot):=\lvert\cdot-b\rvert,b\in[0,1]$. When $F(\cdot,x)$ is a symmetric unimodal distribution with a well-defined conditional density, then $b=1/2$ is optimal in the sense that the interval outputted by Algorithm 1 has minimum length asymptotically [^1] The fact that $b=1/2$ leads to intervals that are asymptotically similar to the oracle has nothing to do with optimal (smallest) length.. An extension to choose $b$ optimally to ensure efficiency as in Chernozhukov, Wüthrich & Zhu (2021) would be certainly possible, but in this article we are interested in LPBs as is natural in survival analysis applications, so that we set $b=1$ unless explicitly stated. Observe that after settng $b=1$,
[…]
As (empirical) survival functions are monotonically decreasing we have that, upon setting $b=1$, the previous set is indeed given by an LPB $(\widehat{L}_{1-\alpha},+\infty)$ such that:
$\widehat{L}_{1-\alpha}(X_{N+1})=\inf\{t:{\hat{S}_{1}\left(t,X_{N+1}\right)\leq\hat{Q}_{\mathcal{I}_{2}}^{*}}\}$
The operating mode $e=0$ is a generalization of Candès et al. (2023) to interval-censoring in virtue of (2), which enjoys finite sample size calibration guarantees but pays the price of overconservativeness; as the original approach does. There is no randomness involved at all when $e=0$.
Theorem 1 (Finite-sample validity of uncervals).
Set $e=0$, or equivalently, just consider the evaluations of $\hat{F}_{1}$ on left borders. Suppose that the data are i.i.d. or exchangeable and that the estimator of the conditional survival function is invariant to permutations of the data. Then $\widehat{\mathcal{C}}_{1-\alpha}^{0}(\cdot)$ is an LPB and
[…]
We study in the next section the theoretical properties of the mode $e=*$, which has provable desirable asymptotic coverage.
2.2 Statistical theory: the interval process
The main goals of this section are to show that uncervals operating with $e=*$ provides asymptotically valid prediction intervals and to derive at which rate this happens. We have introduced uncervals as a meta-algorithm that in practice is constituted by a symbiosis between conformal prediction and the bootstrap. However, from a theoretical perspective its foundations are rooted in a new class of functions whose properties capture the phenomenology of interval data and enable the establishment of performance guarantees through VC theory arguments.
Upon fixing $\hat{F}_{1}$, we have transitioned to an auxiliary sample $(\Lambda_{1},\Upsilon_{1}),(\Lambda_{2},\Upsilon_{2}),\ldots,(\Lambda_{n},\Upsilon_{n})$ which is i.i.d. [or exchangeable if the original sample was so] after having fixed $\hat{F}_{1}$ and respects the constraint $0\leq\Lambda_{i}\leq\Phi_{i}\leq\Upsilon_{i}\leq 1$ P-a.s. for each $1\leq i\leq n$. Classical split distributional conformal prediction relies on the $(1-\alpha)\left(1+1/n\right)$ empirical quantile of $\psi(\Phi_{1}),\ldots,\psi(\Phi_{2})$. In the presence of interval censoring, we do not have access to $\Phi_{1},\ldots,\Phi_{n}$ and even less to the aforementioned empirical quantile. However, we will use the $\Phi_{i}$’s as auxiliary entities to support our theoretical arguments, even though they are never utilized by Algorithm 1.
Remark 1.
Note that simply $\Upsilon_{i}=1$ if observation $i$ is right-censored.
We will use the following conceptual bridge to open a connection between conformal inference and empirical process theory. Suppose $\Phi_{1},\ldots,\Phi_{n}$ are exchangeable random variables free from ties. The rank of the $n$-th observation $\Phi_{n}$ among $\Phi_{1},\ldots,\Phi_{n}$ is uniformly distributed over the set $\{1,\ldots,n\}$. What is to say, under exchangeability we have $\mathbb{P}\left\{\frac{\textrm{rank}(\Phi_{n})}{n}>\epsilon\right\}\geq 1-\epsilon,\textrm{ for all }\epsilon\in[0,1]$. The key connection with empirical process theory is the following observation:
$\frac{\textrm{rank}(\Phi_{n})}{n}$ =#$=\frac{\#\left\{i=1,\ldots,n:\Phi_{i}\geq\Phi_{n}\right\}}{n}$ $=\frac{1}{n}\sum_{i=1}^{n}1\{\Phi_{i}\geq\Phi_{n}\}$ $=1-\texttt{ecdf}_{\Phi_{1},\ldots,\Phi_{n}}(\Phi_{n})$
In this way, the the problem reduces to the study of the asymptotic properties of the empirical distribution function of $\Phi_{1},\ldots,\Phi_{n}$. One of the core mechanisms of split conformal is the following fact: if original data was i.i.d. [resp. exchangeable], then conditional on split 1 scores are i.i.d [resp. exchangeable] even though the model under which $\hat{F}_{1}$ was fit had been misspecified. If we condition on $\mathcal{I}_{1}$, the calibration residuals built by evaluating on split 2 and the test residual are all i.i.d. [resp. exchangeable]. We will see that it is the theoretical mechanism inherent to split conformal of conditioning on a subset of data what makes bootstrap being valid. Under interval censoring, we lose access to the $\Phi_{1},\ldots,\Phi_{n}$, but we will see that it is still possible to recover bootstrap replicates $\Phi^{*}_{1},\ldots,\Phi^{*}_{n}$ from a very particular distribution coined by us the interval measure.
Definition 1.
[…]
Now let $\mathcal{G}=\left\{g_{t}(l,u):t\in[0,1]\right\}$, where
Remark 2.
Remark 2 asserts that when the borders of an inteval collapse into one number, the contribution of such observation to the ecdf is a step function - as happens in the complete information regime.
Let $t\in[0,1]$. Then $\mathbb{I}_{n}(t)=P_{n}g_{t}$. Provided that $\mathbb{E}g_{t}(\Lambda,\Upsilon)=G(t)$, where $G$ is the true cumulative distribution function of the $\Phi_{i}^{\prime}s$, the interval process $\sqrt{n}\left(\mathbb{I}_{n}-G\right)$ coincides with the abstract empirical process indexed by the class $\mathcal{G}$: $\{\sqrt{n}(P_{n}-P)g_{t},\quad g_{t}\in\mathcal{G}\}$. As $\mathcal{G}$ has a finite envelope function, the map $\sqrt{n}(P_{n}-P)=:\mathbb{H}_{n}$ can be viewed as an element in $\ell^{\infty}(\mathcal{G})$ classically known as the empirical process. To summarize:
$\mathbb{Q}_{n}$ Interval measure, $\mathbb{Q}_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\Upsilon_{i}-\Lambda_{i}}1_{(\Lambda_{i},\Upsilon_{i})}(t)$